SINGLE VIBRONIC LEVEL EMISSION SPECTROSCOPY OF CHCl AND CDCl: VIBRATIONAL STRUC-TURE OF THE X^{1}A′ AND a^{3}A′′ STATES

\underline{C. MUKARAKATE}, C. TAO, AND S. A. REID, Department of Chemistry, Marquette University, P.O. Box 1881, Milwaukee, WI 53201-1881.

We report on single vibronic level (SVL) emission spectra from bands in the progressions 2_{0}^{n}, 2_{0}^{n}3_{0}^{m}, 1_{0}^{1}2_{0}^{n}, and 1_{0}^{1}2_{0}^{n}3_{0}^{1} in the A^{1}A′′-X^{1}A′ system of CH^{35}Cl, CH^{37}Cl, CD^{35}Cl and CD^{37}Cl. As in previous studies of CHF/CDF and CHBr/CDBr, the carbenes were generated using a pulsed discharge source, and SVL emission spectra obtained using a 0.3 m spectrograph in combination with a gated, intensified CCD detector. These spectra reveal rich new detail regarding the vibrational structure of the X^{1}A′ and a^{3}A′′ states, and spin-orbit induced mixing between them, up to 9000 cm^{−1} above the vibrationless level of the X^{1}A′ state. For CHCl, we observe around three times the number of X^{1}A′ levels previously reported,^{a}, and a number of new a^{3}A′′ state levels. The results of Dunham expansion fits to the vibrational term energies, and comparisons with previous experimental and recent high quality theoretical studies^{b}^{,}^{c} will be reported. Overall, the derived vibrational parameters of the X^{1}A′ and a^{3}A′′ states are in excellent agreement with ab initio predictions, including our own DFT calculations.

aC.-S. Lin et al., J. Chem. Phys. 121, 4164 (2004).

bG. Tarczay et al., Phys. Chem. Chem. Phys. 7, 2881 (2005).

cH.-G. Yu et al., Mol. Phys. 104, 47 (2006).